1.8g of glucose is added to 1kg of water.what is the molality of solution?

Grade:12

4 Answers

Shailendra Kumar Sharma
6 years ago
molality is moles of solute in per kg of solution.
Now in 1.8g of glucose, no of moles is =1.8/180 =0.01moles of Glucose. 
Total weight of solution 1000+1.8 g =1001.8kg 
So the molality will be .01*1000/1001.8 =.0099 molal
Rohit
6 years ago
The molecular formula of glucose is C6H12O6 it`s molar mass is 180g. Given wt of glucose 1.8g.No. of moles of glucose = 0.01 mol.Wt. of H2O is 1kg. Wt. Of Sol. = 0.0018+1= 1.0018 kg.MOLALITY = moles of solute (glucose)/wt of solution in kg = 0.01/1.0018 = 0.00998 m.
Shriya gupta
6 years ago
Molar weight of glucose is 180 g and the mass given is 1.8 g so mole is equal to 1.8 /180 i. e. Wt/ Molar wt. . It comes out to be 0.01 now molality is mole of solute / wt. Of solvent in kg Hence answer is 0.01/10.01
Prajapati S
5 years ago
Dear Students,
 
Molality is No. of moles of solute present in 1 kg of solvent( not in solution).
Hence Molality = moles of solute/ mass of solvent in kg
= 1.8/180 /1
= 0.01 m
= 1*10-2 m